Can you please help me with these 2 questions?

1). Find an equation in standard form of the parabola described. Vertex (8, 3); focus at (8, 10)

2). Find the required term in the binomial expansion.

(5a − b)8; 7th term
Answer #1

1 ) verten (9.4) = (83) foons (h, K+P) = (8.10) so, h=8, k=3, K+P = 10 -> p = 10 - =z so, Requinea equation of parabola is (nsi uCh-u) to wast 7-(1+aru+/z novo nen h (2-) 8 so, 7th 8 e term of (ha-b) is (5a78-4 (-6) (5ab? (-6) 9 8 = co 70o ab scannea